# Mirror symmetry and rigid structures of generalized K3 surfaces

### v5

@@ 4.3. Mirror symmetry for generalized K3 surfaces
A $(K,L)$-polarized generalized K3 surface is a pair $(X,i)$ of a generalized K3 surface $X=(\varphi,\varphi^{\prime})$ and a primitive embeddings $i:K\oplus L\hookrightarrow H^{*}(M,\mathbb{Z})$ such that $i(K)\subset\widetilde{NS}(X)$ and $i(L)\subset\widetilde{T}(X)$.
Two $(K,L)$-polarized generalized K3 surfaces $(X,i)$ and $(Y,j)$ are called isomorphic if there is a diffeomorphism $f\in\mathrm{Diff}_{*}(M)$ such that $X=f^{*}Y$ and $f^{*}\circ j=i$. For simplicity, we often omit $i$ from the notation $(X,i)$ when there is no confusion.
There may be several families of $(K,L)$-polarized generalized K3 surfaces, and we usually choose one for the purpose of mirror symmetry. This remark is particularly important because of symmetry of Definition 4.5. That is, if $(\varphi,\varphi^{\prime})$ is a $(K,L)$-polarized generalized K3 surface, $(\varphi^{\prime},\varphi)$ is a $(L,K)$-polarized generalized K3 surface in a natural way. Nevertheless, as explained in Section 4.6, we may get a non-trivial result by choosing an appropriate family.
@@
As mentioned in Remark 3.2, in general, the elements in $H^{0}(M,\mathbb{Z})$ and $H^{4}(M,\mathbb{Z})$ are no longer “algebraic” and there is no natural $U\cong H^{0}(M,\mathbb{Z})\oplus H^{4}(M,\mathbb{Z})$ summand in $\widetilde{NS}(X)$. Hence there is no assumption on lattices.
@@
Our formulation is compatible with Aspinwall–Morrison’s discussion from the viewpoint of $N=(2,2)$ SCFT (Section 3 of [2]), where they consider arbitrary $\Omega,\mho\in\widetilde{\mathfrak{D}}$ which are orthogonal, and the mirror map is simply an interchange $\Omega\leftrightarrow\mho$. Furthermore, it is explained that choosing $U$ in $H^{*}(M,\mathbb{Z})$ amounts to choosing a specific $N=(2,2)$ SCFT to represent $N=(4,4)$ SCFT.
@@
By the Torelli type theorem (Theorem 2.10), the construction of moduli spaces works in largely the same way as that for lattice polarized K3 surfaces, and hence we will give a brief sketch below (we refer the reader to Dolgachev’s article [3] for details). We first define
$\mathfrak{D}_{K}=\{\varphi\in\mathbb{P}(K_{\mathbb{C}})\ |\ \langle\varphi,\varphi\rangle=0,\langle\varphi,\overline{\varphi}\rangle>0\}\subset\widetilde{\mathfrak{D}}.$
It is a complex manifold of dimension $\kappa-2$ with two connected components $\mathfrak{D}_{K}^{+}$ and $\mathfrak{D}_{K}^{-}$, each of which is a symmetric domain of type IV isomorphic to $\mathrm{Gr}^{po}_{2}(K_{\mathbb{R}})\cong\mathrm{O}(2,\kappa-2)/(\mathrm{SO}(2)\times\mathrm{O}(\kappa-2))$. By a slight modification of the Torelli type theorem, the $A$-moduli space $\mathfrak{M}_{A}$ is identified with an open dense subset of the period domain $\mathfrak{D}_{K}^{+}$. The construction is parallel for the $B$-moduli space $\mathfrak{M}_{B}$. Therefore, the full moduli space $\mathfrak{M}_{A}\times\mathfrak{M}_{B}$ of $(K,L)$-polarized generalized K3 surfaces is always $20$-dimensional and identified with an open dense subset of $\mathfrak{D}_{(K,L)}=\mathfrak{D}_{K}^{+}\times\mathfrak{D}_{L}^{+}$.
@@ 4.4. Comparison: classical and new
We will next confirm that our formulation of mirror symmetry naturally includes Dolgachev’s.
Let $K^{\prime}\subset\Lambda_{K3}$ be a sublattice of signature $(1,\rho-1)$. Assume that there exists a decomposition $K^{\prime\perp}=L^{\prime}\oplus U$ for a lattice $L^{\prime}$. Mirror symmetry for lattice polarized K3 surfaces asserts that the $K^{\prime}$-polarized K3 surfaces and the $L^{\prime}$-polarized K3 surfaces are mirror symmetric to each other.
Let $M_{\sigma}$ be a generic $K^{\prime}$-polarized K3 surface equipped with a $B$-field $B$ shift of a hyperKähler form $\omega$ ($B,\omega\in NS(M_{\sigma})_{\mathbb{R}}$ are also taken generically). It defines a generalized K3 surface $X=(\varphi,\varphi^{\prime})=(e^{B+\sqrt{-1}\omega},\sigma)$ and we have
@@
Similarly, let $M_{\sigma^{\vee}}$ be a generic $L^{\prime}$-polarized K3 surface equipped with a $B$-field $B^{\vee}$ shift of a hyperKähler form $\omega^{\vee}$. It defines a generalized K3 surface $X^{\vee}=(\varphi,\varphi^{\prime})=(e^{B^{\vee}+\sqrt{-1}\omega^{\vee}},\sigma^{\vee})$. and we have
$\widetilde{NS}(X^{\vee})=NS^{\prime}(M_{\sigma^{\vee}})\cong L^{\prime}\oplus U,\ \ \ \widetilde{T}(X^{\vee})=T(M_{\sigma^{\vee}})\cong K\oplus U.$
By setting $K=K^{\prime}\oplus U$ and $L=L^{\prime}\oplus U$, we may think that $X$ is a $(K,L)$-polarized generalized K3 surface and $X^{\vee}$ is a $(L,K)$-polarized generalized K3 surface. Therefore the formulation of mirror symmetry for lattice polarized K3 surfaces is naturally considered as a special case of ours.
4.5. Mirror symmetry and $B$-fields
Our formulation of mirror symmetry comes into its own when the conventional formulation fails. Let us consider the first problem in Section 4.2.
Let $K^{\prime}\subset\Lambda_{K3}$ be a sublattice of signature $(1,\rho-1)$. Assume that there exists a decomposition $K^{\prime\perp}=L^{\prime}\oplus U(m)$ for a lattice $L^{\prime}$ and $m\geq 2$. We set $K=K^{\prime}\oplus U(m)$ and $L=L^{\prime}\oplus U(m)$.
A $(K,L)$-polarized generalized K3 surface can be constructed as follows. First we consider $K^{\prime}$-polarized K3 surface $M_{\sigma}$ so that generically $T(M_{\sigma})\cong L^{\prime}\oplus U(m)$. Let $e,f$ be the standard basis of this $U(m)$, i.e. $e^{2}=f^{2}=0$ and $(e,f)=m$, and consider a $B$-field $B^{\prime}=\frac{1}{m}e\in T(M_{\sigma})_{\mathbb{Q}}$. Then, for generic $B,\omega\in NS(M_{\sigma})_{\mathbb{R}}$,
@@ 4.5. Mirror symmetry and $B$-fields
is a $(K,L)$-polarized generalized K3 surface. In fact, by the calculation in Example 3.3,
$\widetilde{NS}(X)=NS(M_{\sigma})\oplus\mathbb{Z}(m,e,0)\oplus H^{4}(M_{\sigma},\mathbb{Z})\cong K^{\prime}\oplus U(m)$
@@
$\widetilde{T}(X)=e^{B^{\prime}}T(M_{\sigma})_{\mathbb{Q}}\cap H^{*}(M_{\sigma},\mathbb{Z})\cong T(M_{\sigma})\cong L^{\prime}\oplus U(m).$
An $(L,K)$-polarized K3 surface is defined in the same manner. Note that the elements in $H^{0}(M,\mathbb{Z})$ are no longer algebraic in this case.
A $B$-field transform of a complex structure $\sigma$ is thought of as a non-commutative deformation of a K3 surface $M_{\sigma}$ in the sense that the geometric objects are twisted by the Brauer group $\alpha$ associated to the $B$-field. The corresponding category is the derived category of twisted coherent sheaves $\mathrm{D}^{b}(M_{\sigma},\alpha)$ (cf. Section 4.7).
@@ 4.6. Mirror symmetry for singular K3 surfaces
Let us next consider the second problem in Section 4.2, namely mirror symmetry for singular (i.e. complex rigid) K3 surfaces.
@@
$K=\langle-2n\rangle^{\oplus 2}\oplus U\oplus E_{8}^{\oplus 2},\ \ \ L=\langle 2n\rangle^{\oplus 2},$
where $\langle k\rangle$ denotes the lattice of rank $1$ generated by $v$ with $v^{2}=k$.
@@
    A family $\mathcal{X}$ of the $(K,L)$-polarized generalized K3 surfaces is given by $\{X=(e^{B+\sqrt{-1}\omega},\sigma)\}$, where $T(M_{\sigma})=L$ (such a complex structure is unique up to isomorphism) and $B,\omega\in NS(M_{\sigma})_{\mathbb{R}}$. In other words, they are singular K3 surfaces with complexified Kähler parameters $B+\sqrt{-1}\omega\in NS(M_{\sigma})_{\mathbb{C}}$. For a generic $X$, we have
@@
    A family $\mathcal{Y}$ of the $(L,K)$-polarized generalized K3 surfaces contains a 19-dimensional subfamily of the K3 surfaces of the form $\{Y=(e^{\sqrt{-1}H},\sigma^{\vee})\}$, where $NS(M_{\sigma^{\vee}})=\mathbb{Z}H,H^{2}=2n>0$. There is no Kähler deformation (cf. Section 3.4). In fact, for a generic $X^{\vee}$, we have
@@
We observe that the dimensions of the $A$- and $B$-moduli spaces are interchanged between the mirror families.
@@
The discussion above is readily generalized to any singular K3 surfaces. Let $S$ be a singular K3 surface such that $NS(S)=K^{\prime}$ and $T(S)=L$, and set $K=K^{\prime}\oplus U$. Then the mirror is a family $\mathcal{Y}$ of $(L,K)$-polarized generalized K3 surfaces. This family may contain a K3 surface with $L_{e^{B+\sqrt{-1}\omega}}\cong L$, but an explicit construction of such a K3 surface is normally very hard.
To summarize, the mirror partner of a singular K3 surfaces is a Kähler rigid generalized K3 surface, which is in general not a classical K3 surface. This solves the long-standing problem of mirror symmetry for singular K3 surfaces.
@@
The initial motivation of the present article comes from the attractor mechanisms of the complex and Kähler moduli spaces of Calabi–Yau 3-folds [4]. The connection is given by a Calabi–Yau 3-fold defined by the product of an elliptic curve and a K3 surface. Then the complex (resp. Kähler) attractor mechanism specifies a set of isolated points in the moduli space, called the constellation, and they correspond to the complex (resp. symplectic) rigid K3 surfaces. See [4] for details.

### v6

@@ 4.3. Mirror symmetry for generalized K3 surfaces
A $(K,L)$-polarized generalized K3 surface is a pair $(X,i)$ of a generalized K3 surface $X=(\varphi,\varphi^{\prime})$ and a lattice embedding $i:K\oplus L\hookrightarrow H^{*}(M,\mathbb{Z})$ such that
  1. (1)
    the restrictions $i|_{K}$ and $i|_{L}$ are primitive embeddings,
  2. (2)
    $i(K)\subset\widetilde{NS}(X)$ and $i(L)\subset\widetilde{T}(X)$.
Two $(K,L)$-polarized generalized K3 surfaces $(X,i)$ and $(Y,j)$ are called isomorphic if there is a diffeomorphism $f\in\mathrm{Diff}_{*}(M)$ such that $X=f^{*}Y$ and $f^{*}\circ j=i$. For simplicity, we usually omit $i$ from the notation $(X,i)$ when there is no confusion.
@@
As mentioned in Remark 3.2, in general, the elements in $H^{0}(M,\mathbb{Z})$ and $H^{4}(M,\mathbb{Z})$ are no longer “algebraic” and there is no natural summand $U\cong H^{0}(M,\mathbb{Z})\oplus H^{4}(M,\mathbb{Z})$ in $\widetilde{NS}(X)$. Therefore, there is no assumption on the lattices $K$ and $L$ in our formulation.
There may be several families of $(K,L)$-polarized generalized K3 surfaces, and we typically select one for the purpose of mirror symmetry. This remark is particularly important due to the symmetry of Definition 4.4. Specifically, if $(\varphi,\varphi^{\prime})$ is a $(K,L)$-polarized generalized K3 surface, then $(\varphi^{\prime},\varphi)$ is a $(L,K)$-polarized generalized K3 surface in a natural way. Nevertheless, as elucidated in Section 4.5 and Section 4.6, we may obtain a non-trivial result by selecting appropriate families. This issue is also related to the so-called multiple mirrors phenomenon.
@@
Our formulation is compatible with Aspinwall–Morrison’s description of mirror symmetry for K3 surfaces from the viewpoint of SCFTs [3]. They identified the moduli space of $N=(2,2)$ SCFTs on a K3 surface with the Grassmannian
$\mathrm{Gr}_{2,2}^{po}(H^{*}(M,\mathbb{R}))\cong\mathrm{O}(4,20)/(\mathrm{SO}(2)\times\mathrm{SO}(2)\times\mathrm{O}(20))$
parametrizing the orthogonal pairs of positive oriented 2-planes in $H^{*}(M,\mathbb{R})\cong\mathbb{R}^{4,20}$. To a K3 surface $M_{\sigma}$ with a Kähler class $\omega$, we may associate
$(P_{\sigma},P_{e^{\sqrt{-1}\omega}})\in\mathrm{Gr}_{2,2}^{po}(H^{*}(M,\mathbb{R})).$
There are “non-geometric points” that cannot be obtained in this way. However, we now understand that these points are given by generalized K3 surfaces, as stated in Theorem 2.11.
Then mirror symmetry is an involution of $\mathrm{Gr}_{2,2}^{po}(H^{*}(M,\mathbb{R}))$ given by swapping the two 2-planes, which does not preserve the geometric points. Therefore, it is understood that a mirror partner of a K3 surface is not necessarily a K3 surface, but rather a generalized K3 surface. It is also explained that choosing $U$ in $H^{*}(M,\mathbb{Z})$ amounts to selecting a specific $N=(2,2)$ SCFT to represent an $N=(4,4)$ SCFT.
@@
We will provide an explicit description of the moduli space $\mathfrak{M}_{A}\times\mathfrak{M}_{B}$ of the $(K,L)$-polarized generalized K3 surfaces. We begin by introducing
$\mathfrak{D}_{K}:=\{\varphi\in\mathbb{P}(K_{\mathbb{C}})\ |\ \langle\varphi,\varphi\rangle=0,\langle\varphi,\overline{\varphi}\rangle>0\},$
a complex manifold of dimension $\kappa-2$, isomorphic to the Grassmannian $\mathrm{Gr}^{po}_{2}(K_{\mathbb{R}})\cong\mathrm{O}(2,\kappa-2)/(\mathrm{SO}(2)\times\mathrm{O}(\kappa-2))$ parametrizing the positive oriented 2-planes in $K_{\mathbb{R}}\cong\mathbb{R}^{2,\kappa-2}$. It has two connected components, each of which is isomorphic to a symmetric domain of type IV. $\mathfrak{D}_{L}$ is defined in a similar manner.
Theorem 4.8.
The moduli space $\mathfrak{M}_{A}\times\mathfrak{M}_{B}$ of the $(K,L)$-polarized generalized K3 surfaces is identified with an open dense subset of $\mathfrak{D}_{(K,L)}:=\mathfrak{D}_{K}\times\mathfrak{D}_{L}$. Moreover, $\mathfrak{M}_{A}$ is identified with an open dense subset of $\mathfrak{D}_{K}$, and similarly, $\mathfrak{M}_{B}$ is identified with an open dense subset of $\mathfrak{D}_{K}$.
Proof.
Let $(X=(\varphi,\varphi^{\prime}),i)$ be a $(K,L)$-polarized generalized K3 surface. Then the conditions $i(K)\subset\widetilde{NS}(X)$ and $i(L)\subset\widetilde{T}(X)$ are equivalent to $\varphi^{\prime}\in\mathfrak{D}_{i(L)}$ and $\varphi\in\mathfrak{D}_{i(K)}$, respectively. We will demonstrate that a point in an open dense subset of $\mathfrak{D}_{i(K)}\times\mathfrak{D}_{i(L)}$ defines a unique generalized K3 surface. Firstly, we associate to $(\alpha,\beta)\in\mathfrak{D}_{i(K)}\times\mathfrak{D}_{i(L)}$ an orthogonal pair of positive oriented 2-planes $(P_{\alpha},P_{\beta})\in\mathrm{Gr}^{po}_{2,2}(H^{*}(M,\mathbb{R}))$, which are generically not orthogonal to any $(-2)$-class. Therefore, by Theorem 2.11, the period map $\mathfrak{per}_{\mathrm{gK3}}:\mathfrak{M}_{\mathrm{gK3}}\rightarrow\mathrm{Gr}^{po}_{2,2}(H^{*}(M,\mathbb{R}))$ ensures that, generically, $(\alpha,\beta)\in\mathfrak{D}_{i(K)}\times\mathfrak{D}_{i(L)}$ defines a unique generalized K3 surface. Consequently, the moduli space $\mathfrak{M}_{A}\times\mathfrak{M}_{B}$ may be identified with an open dense subset of $\mathfrak{D}_{i(K)}\times\mathfrak{D}_{i(L)}\cong\mathfrak{D}_{(K,L)}$.
Let $H:=(H^{2}(M,\mathbb{C})\oplus H^{4}(M,\mathbb{Z}))\cap i(K)$. Then, by Theorem 2.12 and the above discussion, $\mathfrak{D}_{i(K)}\setminus\mathbb{P}(H)$ parametrizes the $A$-deformation of $(\varphi,\varphi^{\prime})$ for a generic choice of $\varphi^{\prime}\in\mathfrak{D}_{i(L)}$. Therefore, $\mathfrak{M}_{A}$ is identified with $\mathfrak{D}_{i(K)}\setminus\mathbb{P}(H)$. A parallel discussion can be carried out for $\mathfrak{M}_{B}$. ∎
@@ 4.4. Comparison: classical and new
Next, we will demonstrate that our formulation of mirror symmetry inherently encompasses Dolgachev’s formulation.
For an integer $1\leq\rho\leq 19$, let $K^{\prime}$ be a lattice of signature $(1,\rho-1)$. Assume that there exists a decomposition $K^{\prime\perp}=L^{\prime}\oplus U$ for a lattice $L^{\prime}$. Mirror symmetry for lattice polarized K3 surfaces posits that the $K^{\prime}$-polarized K3 surfaces and the $L^{\prime}$-polarized K3 surfaces are mirror symmetric to each other.
Let $M_{\sigma}$ be a generic $K^{\prime}$-polarized K3 surface equipped with a $B$-field $B$ shift of a hyperKähler form $\omega$ ($B,\omega\in NS(M_{\sigma})_{\mathbb{R}}$ are also taken generically). It defines a generalized K3 surface $X=(\varphi,\varphi^{\prime})=(e^{B+\sqrt{-1}\omega},\sigma)$. By the calculation in Example 3.3,
@@
Similarly, let $M_{\sigma^{\vee}}$ be a generic $L^{\prime}$-polarized K3 surface equipped with a $B$-field $B^{\vee}$ shift of a hyperKähler form $\omega^{\vee}$. It defines a generalized K3 surface $X^{\vee}=(\varphi,\varphi^{\prime})=(e^{B^{\vee}+\sqrt{-1}\omega^{\vee}},\sigma^{\vee})$ and
$\widetilde{NS}(X^{\vee})=NS^{\prime}(M_{\sigma^{\vee}})\cong L^{\prime}\oplus U,\ \ \ \widetilde{T}(X^{\vee})=T(M_{\sigma^{\vee}})\cong K^{\prime}\oplus U.$
By setting $K:=K^{\prime}\oplus U$ and $L:=L^{\prime}\oplus U$, we may think that $X$ is a $(K,L)$-polarized generalized K3 surface and $X^{\vee}$ is an $(L,K)$-polarized generalized K3 surface. Therefore the formulation of mirror symmetry for lattice polarized K3 surfaces is naturally regarded as a special case of ours.
4.5. Mirror symmetry and $B$-field twists
Our formulation of mirror symmetry comes into its own when the conventional approach falls short. Let us delve into the first issue outlined in Section 4.2.
For an integer $1\leq\rho\leq 19$, let $K^{\prime}$ be a lattice of signature $(1,\rho-1)$. Assume that there exists a decomposition $K^{\prime\perp}=L^{\prime}\oplus U(m)$ for a lattice $L^{\prime}$ and $m\geq 2$. We define $K:=K^{\prime}\oplus U(m)$ and $L:=L^{\prime}\oplus U(m)$.
A $(K,L)$-polarized generalized K3 surface can be constructed as follows. First we consider $K^{\prime}$-polarized K3 surface $M_{\sigma}$ so that generically $T(M_{\sigma})\cong L^{\prime}\oplus U(m)$. Let $e,f$ be the standard basis of this $U(m)$, i.e. $e^{2}=f^{2}=0$ and $\langle e,f\rangle=m$, and consider a $B$-field $B^{\prime}=\frac{1}{m}e\in T(M_{\sigma})_{\mathbb{Q}}$. Then, for $B,\omega\in NS(M_{\sigma})_{\mathbb{R}}$,
@@ 4.5. Mirror symmetry and $B$-field twists
is a $(K,L)$-polarized generalized K3 surface. Indeed, for generic $X$, by the calculation in Example 3.3,
$\widetilde{NS}(X)$ $=NS(M_{\sigma})\oplus\mathbb{Z}(m,e,0)\oplus H^{4}(M_{\sigma},\mathbb{Z})$
$\cong K^{\prime}\oplus U(m)$
@@
$\widetilde{T}(X)$ $=e^{B^{\prime}}T(M_{\sigma})_{\mathbb{R}}\cap H^{*}(M_{\sigma},\mathbb{Z})$
$=L^{\prime}\oplus\mathbb{Z}(0,e,0)\oplus\mathbb{Z}(0,f,1)$
$\cong L^{\prime}\oplus U(m).$
Note that the elements in $H^{0}(M,\mathbb{Z})$ are no longer algebraic in this case. An $(L,K)$-polarized K3 surface can be constructed in the same manner.
Remark 4.9.
A $B$-field shift of a complex structure $\sigma$ is thought of as a non-commutative deformation of a K3 surface $M_{\sigma}$ where the geometric objects are twisted by the Brauer group element $\alpha$ associated with the $B$-field. The corresponding category is the derived category of twisted coherent sheaves $\mathrm{D}^{b}(M_{\sigma},\alpha)$. See for example [13].
@@ 4.6. Mirror symmetry for singular K3 surfaces
Let us next consider the second problem in Section 4.2, namely the problem of mirror symmetry for singular K3 surfaces.
@@
$K:=\langle-2n\rangle^{\oplus 2}\oplus U^{\oplus 2}\oplus E_{8}^{\oplus 2},\ \ \ L:=\langle 2n\rangle^{\oplus 2}.$
@@
    A family $\mathcal{X}$ of the $(K,L)$-polarized generalized K3 surfaces is given by $\{X=(e^{B+\sqrt{-1}\omega},\sigma)\}$, where $T(M_{\sigma})=L$ and $B,\omega\in NS(M_{\sigma})_{\mathbb{R}}$. Note that $M_{\sigma}$ is realized as a $(\langle-2n\rangle^{\oplus 2}\oplus U\oplus E_{8}^{\oplus 2})$-polarized K3 surface. For generic $X$, we have
@@
    A family $\mathcal{Y}$ of the $(L,K)$-polarized generalized K3 surfaces contains a 19-dimensional subfamily of the K3 surfaces of the form $\{Y=(e^{\sqrt{-1}H},\sigma^{\vee})\}$, where $NS(M_{\sigma^{\vee}})=\mathbb{Z}H$ with $H^{2}=2n>0$. There is no Kähler deformation as explained in Section 3.4. Indeed, for generic $X^{\vee}$, we have
@@
We therefore obtain two $20$-dimensional families of generalized K3 surfaces whose $A$-moduli space and $B$-moduli space are exchanged as expected (cf. Theorem 4.8).
@@
The above discussion is readily generalized to any singular K3 surfaces. Let $S$ be a singular K3 surface such that $NS(S)=K^{\prime}$ and $T(S)=L$, and set $K=K^{\prime}\oplus U$. Then the mirror is given by a family $\mathcal{Y}$ of $(L,K)$-polarized generalized K3 surfaces. This family may include a K3 surface with $L_{e^{B+\sqrt{-1}\omega}}\cong L$, although it is normally very hard to construct such a K3 surface explicitly.
In summary, the mirror partner of a singular K3 surfaces is a Kähler rigid generalized K3 surface, which is generally not a classical K3 surface. This solution addresses the problem of mirror symmetry for singular K3 surfaces.
@@
The initial motivation for the present article stems from the attractor mechanisms of the complex and Kähler moduli spaces of Calabi–Yau 3-folds [9]. The connection arises from a Calabi–Yau 3-fold formed by the product of an elliptic curve and a K3 surface. Then the complex (resp. Kähler) attractor mechanism specifies a set of isolated points in the moduli space, called the constellation, and they correspond to the complex (resp. symplectic) rigid K3 surfaces. For further elaboration, refer to [9].